OpenStudy (anonymous):

You roll two dice. What is the probability of rolling a 2 or a 4 on the second die, given that you rolled an even number on the first die? A 6X6 table of dice outcomes will help you to answer this question. A.2/36 B.1/2 C.7/36 D.1/6 E.1/3

OpenStudy (anonymous):

1/3?

OpenStudy (shubhamsrg):

1/6 no ?

OpenStudy (anonymous):

Let A be the event where an even number is rolled on first die Let B be the event where a 2 or 4 is rolled on the 2nd die This is a conditional probability, condition given is to roll an even number on the first die. So we need to find P(BlA). P(BlA) = P(B intersect A) / P(A) P(A) = 3/6=1/2 ( 3 possibilities for even number 2,4,6) P(B intersect A) can be found from the 6x6 table. so the outcomes which satisfy an even number on the first die and 2 or 4 on the second die would be (2,2) (2,4) (4,4) (4,2) (6,2) (6,4). total outcomes would be 6x6=36 so the probability of P(B intersect A) = 6 /36 = 1/6 so P(BlA) = (1/6) / (1/2) = 1/3

OpenStudy (anonymous):

yeah should be 1/3 , as the probability of getting 2 or 4 from throwing a dice is constantly 2 out of 6, and is independent of other throws.

OpenStudy (shubhamsrg):

yes,,sorry..